# DIRESA, a distance-preserving nonlinear dimension reduction technique based on regularized autoencoders

### v1

@@ Abstract
In meteorology, finding similar weather patterns or analogs in historical datasets can be useful for data assimilation, forecasting, and postprocessing. In climate science, analogs in historical and climate projection data are used for attribution and impact studies. However, most of the time, those large weather and climate datasets are nearline. They must be downloaded, which takes a lot of bandwidth and disk space, before the computationally expensive search can be executed. We propose a dimension reduction technique based on autoencoder (AE) neural networks to compress those datasets and perform the search in an interpretable, compressed latent space. A distance-regularized Siamese twin autoencoder (DIRESA) architecture is designed to preserve distance in latent space while capturing the nonlinearities in the datasets. Using conceptual climate models of different complexities, we show that the latent components thus obtained provide physical insight into the dominant modes of variability in the system. Compressing datasets with DIRESA reduces the online storage and keeps the latent components uncorrelated, while the distance (ordering) preservation and reconstruction fidelity robustly outperform Principal Component Analysis (PCA) and other dimension reduction techniques such as UMAP or variational autoencoders.
@@ 1 Introduction
Today, weather and climate models and observation systems generate unprecedented volumes of data. For example, the Coupled Model Intercomparison Project phase 6 (CMIP6) represents more than 20 PB of climate model output data [li2023big], and the world’s Earth Observation data collections grow by around 100 PB per year [wilkinson2024].
@@
These unwieldy archives of Earth system data provide enormous opportunities as well as challenges. They are usually stored in formats such as NetCDF, which has natively supported zlib-based compression since version 4.0. This compression technique does not consider the specific context of climate data, and as a result, many more bits of information are stored than are meaningful or necessary [klower2021compressing].
@@
To allow for the full potential of this vast treasure of data to be unlocked, an approach is needed to reduce these datasets into a domain-aware and physically meaningful space. The approach presented in this work uses a Siamese twin AE with regularization terms to anchor the latent space to meaningful properties such as distance (ordering) preservation and statistical independence.
@@
A major application that benefits from long records of reanalysis and climate model data is the search for analogs. Methods based on analogs, which rely on the retrieval of similar weather patterns, are used for forecasting [vandendool1989new], as Lorenz already proposed in 1969 [lorenz1969atmospheric], or for the calculation of the local attractor properties, such as the instantaneous dimension, a proxy for the predictability [faranda2017dynamical]. In analog data assimilation, the analogs can be searched in historical data archives [lguensat2017analog] or constructed via generative machine learning methods, such as variational AE [yang2021machine] or generative adversarial networks. Besides statistical and dynamical downscaling, higher resolution weather and climate information can be obtained using analog methods [zorita1999analog, ghilain2022large] or neural network-based techniques [rozoff2022comparison]. Finally, analogs are a cornerstone of a novel climate change attribution approach that accounts for changes in atmospheric circulation patterns [faranda2022climate].
@@
As a downside to the growing datasets, however, looking up analogs in these vast records becomes increasingly computationally expensive as the dimensionality of the data and length of the dataset increases. As looking up analogs by minimizing Euclidian distance can be practically infeasible, many approaches have been proposed to quickly retrieve similar patterns of one or more atmospheric variables. These are based on dimensionality reduction (DR) techniques such as the popular classical method PCA (Principal Component Analysis). To capture the nonlinearity of the data, other techniques were developed [van2009dimensionality]. UMAP (Uniform Manifold Approximation and Projection) is one such technique [mcinnes2018umap], which can be combined with advanced search algorithms to find the analog in a fast and accurate way [franch2019mass]. DR is also combined with other data analysis techniques, such as clustering [neal2016flexible] [chattopadhyay2020predicting], or topological techniques, such as persistent homology, for finding weather regimes.
@@
Autoencoders are another technique, based on deep learning, to produce a lower-dimensional latent representation [manning2018pca]. However, they lack the statistical independence of the latent space components and the ordering of these components. The latter can be overcome by adding a MaskLayer [royen2021masklayer]. Decorrelation of the components can be achieved by adding a covariance regularization term on the latent space layer [ladjal2019pca].
@@
When using autoencoders, the Euclidian distance between two samples in the original space is not necessarily correlated with the distance in latent space. Our architecture uses a Siamese twin AE (see Figure 1) to preserve distance (ordering) in latent space. The two encoder branches share the weights and produce two different latent space representations. These are used for the distance loss term, similar to the invariance term used in [bardes2021vicreg], which ensures that the distance between samples in the dataset is reflected (preserved or correlated) in the distance between latent representations of those samples.
@@ 2.1 Lorenz ’63
The Lorenz ’63 system [lorenz1963deterministic] is a 3-dimensional simplified model for atmospheric convection. This system of three ordinary differential equations (ODE) displays chaos and features a strange attractor with the famous butterfly-wings shape for specific ranges of its parameters [palmer1993extended]. The 3-dimensional Lorenz butterfly is compressed into a 2-dimensional latent space by the different DR techniques. The system of ODEs with parameters $\sigma=10$, $r=28$, and $b=8/3$ was integrated using the 4th-order Runge-Kutta numerical scheme, from starting point (1,0,1) with a timestep of 0.0025. The first 1000 points are ignored as they are considered part of the transient. The next 100000 steps are split into a training dataset (first 80000), a validation (next 10000), and a test dataset (last 10000). Before feeding the data into the DR methods, it is scaled to $[0,1]$. Important to remark is that the Hausdorff fractal dimension of the Lorenz butterfly is about 2.06, so it is theoretically impossible to map the 3-dimensional dataset into two dimensions while retaining all the information.
@@ 2.2 Modular Arbitrary-Order Ocean-Atmosphere Model
MAOOAM (Modular Arbitrary-Order Ocean-Atmosphere Model) is a quasigeostrophic coupled ocean-atmosphere model for midlatitudes [de2016modular, VDDG2015]. It features a 2-layer atmosphere that is mechanically and thermodynamically coupled to a shallow-water ocean layer in the $\beta$-plane (linearized Coriolis) approximation. The DDV2016 setup was used, with an atmospheric and oceanic resolution of 2x–2y and 2x–4x, resulting in a 36-dimensional model, with ten wavenumbers for the atmospheric stream function and temperature and eight for the ocean stream function and temperature. The dataset was obtained by integrating the model with the 2nd-order Runge-Kutta numerical scheme (with nondimensional timestep 0.01) for $84e$ 8 timesteps, with a write-out every 100th timestep (1 timestep is $1/f_{0}$ s, with $f_{0}$ the Coriolis parameter at 45°N, which is about 2.7 hours). The first 1M points ($10^{8}$ timesteps) are ignored as part of the transient, required due to the slow oceanic dynamics. For the following 3M points, which represent 92000 years of data, the wavenumbers are converted to a grid of 8 x 6 and four channels (for the four variables) and then split up into a training (first 2.4M), a validation (next 300K), and a test dataset (last 300K).
@@ 3.1 PCA and KPCA
PCA linearly transforms the data into new coordinates, the principal components [pearson1901liii, hotelling1933analysis]. These components are the eigenvectors of the covariance matrix of the dataset. The amount of dataset variation in the direction of each eigenvector determines the ordering of the coordinates. As the covariance matrix is symmetric, the eigenvalues are real, and the eigenvectors are orthogonal. For dimensional reduction, the first principal components are taken as they explain most of the variance in the dataset.
In KPCA, the traditional PCA is combined with kernel methods [scholkopf1997kernel]. The original data points are mapped non-linearly on a high dimensional Hilbert space, where the PCA is applied. The so-called kernel trick is used to avoid working in this higher dimension (and so being highly computationally expensive). The calculations are done in the original space, replacing the standard inner product with a kernel function representing the inner product in the Hilbert space.
@@ 3.2 UMAP
UMAP is based on manifold learning and topological data analysis [mcinnes2018umap]. It creates a fuzzy topological representation of the data, a weighted graph, by connecting the points within a certain radius, which is calculated locally based on the distance to the nearest neighbors. A stochastic gradient descent algorithm with a cross-entropy loss function searches for a low-dimensional representation with the closest possible similar fuzzy topological structure.
@@ 3.3 Autoencoders
An AE is a deep-learning ANN that consists of two parts: an encoder that compresses the data into a latent representation and a decoder that decompresses this representation into its original form [kramer1991nonlinear]. The AE is trained by minimizing the reconstruction loss, which ensures that the output of the decoder is as similar as possible to the encoder’s input. We use the MSE between the encoder’s input and the decoder’s output as reconstruction loss, which can also be seen as a measure for the unexplained variance of the latent representation. After the training and testing of the neural network are done, the original dataset can be compressed with the encoder.
@@
In a VAE, the latent components are not single numbers but probability distributions, usually Gaussian distributions [kingma2022autoencoding]. The encoder has three output layers: one for the means, one for the variances (or logarithm of the variances) of the distributions, and one random sampling taken from those distributions that feed into the decoder submodel. A VAE is a generative method, as each sample of the latent distributions produces a different output after decoding. A VAE has two loss functions. The first one is the reconstruction loss, as in a standard AE. The second loss function contains a Kullback–Leibler divergence regularization term (KL loss), which forces the latent distributions to be standard normal distributions. The total loss function is the weighted average of both losses.
Figure 2: Requirements for the autoencoder. (a) Reconstructed data as close as possible to the original. (b) Distance in original space is related to distance in latent space. (c) Independence of latent components and ordering of the latent components in terms of importance.
@@ 3.4 DIRESA
    Reconstructed dataset as close as possible to the original dataset
@@
    The distance between points in the original space is reflected in their distance in latent space
@@
    The latent components are independent of each other
@@
    The latent components are ordered in terms of importance
@@
The DIRESA architecture (Figure 1) fulfils the above requirements. As in a standard AE, the reconstruction loss enforces the first requirement. We use the MSE between the original data, fed into the encoder, and the decoder output as reconstruction loss.
@@
A shuffled version of the dataset is fed into the twin encoder, which shares the weights and biases with the original encoder. The random ordering of the shuffled dataset can be done beforehand on the complete dataset (and so does not change during training) or can be executed per batch during training by the data generator. We used the first option here to be independent of the batch size. A distance layer calculates the Euclidian distance $d(x,x^{\prime})$ between the original and shuffled input and between the latent components of the two twins, $d(z,z^{\prime})$. Minimizing the distance loss forces the relationship between those two. For DIRESA, several distance loss functions have been implemented. DIRESAMSE uses the MSE between $d(x,x^{\prime})$ and $d(z,z^{\prime})$, meaning that distance is preserved between original and latent space. The MSLE, which calculates the MSE between $log(d(x,x^{\prime})+1)$ and $log(d(z,z^{\prime})+1)$, is used for DIRESAMSLE, focusing more on short than long distances. DIRESACorr uses a distance Pearson correlation loss: $Loss_{Corr}=1-corr(d(x,x^{\prime}),d(z,z^{\prime}))$, forcing the correlation between original and latent space distances. The distance Pearson correlation loss is calculated per batch (one figure per batch and not one figure per sample as in an MSE loss function), so the batch size must be large enough to have an accurate figure. A last distance loss function is used in DIRESALogCorr, where the correlation loss is computed on the logarithm of the distances plus one.
@@
The independence of the latent components is forced by the covariance loss $Loss_{Cov}=\sum_{i\neq j}cov^{2}_{i,j}(z)/(L*(L-1))$ which is the normalized squared sum of the non-diagonal terms of the covariance matrix, with $L$ the number of dimensions of the latent space. The denominator keeps the loss function independent of the latent space dimension. The covariance matrix is calculated by batch, so again, it is essential to have the batch size big enough so that the batch covariance is a good approximation of the dataset covariance.
@@
The total loss is a weighted average of the three different loss components. For lowering the weight factor tuning effort, annealing [behrens2022non] is used for the covariance loss, meaning that the weight factor is 0 at the start and gradually increases. Annealing stops when the covariance loss reaches a point, where we consider the latent components independent. The covariance regularized AE (CRAE) is similar to DIRESA but with only the covariance loss and no distance loss (and so no twin encoder).
@@
The last requirement, the ordering of the latent components, could be imposed by a Masklayer, but in practice, it was found to hamper the training of the neural network. So, the ordering is not imposed during training but is calculated afterwards. To achieve this, all encoded (latent) dataset elements are set to the average value except for one latent component. After repeating this process for all latent components and decoding them, the latent components can be ordered based on their variances in the decoded (original) space.
@@
Figure 3: Test dataset (blue), next to the decoded random uniform latent components (orange) for PCA (3(a)), KPCA (3(b)), UMAP (3(c)), AE (3(d)), BNAE (3(e)), CRAE (3(f)), VAE (3(g)), DIRESAMSE (3(h)) and DIRESACorr (3(i))
Figure 4: Encoded Lorenz test dataset for PCA (4(a)), KPCA (4(b)), UMAP (4(c)), AE (4(d)), BNAE (4(e)), CRAE (4(f)), VAE (4(g)), DIRESAMSE (4(h)) and DIRESACorr (4(i))

### v2

@@ Abstract
In meteorology, finding similar weather patterns or analogs in historical datasets can be useful for data assimilation, forecasting, and postprocessing. In climate science, analogs in historical and climate projection data are used for attribution and impact studies. However, most of the time, those large weather and climate datasets are nearline. This means that they must be downloaded, which takes a lot of bandwidth and disk space, before the computationally expensive search can be executed. We propose a dimension reduction technique based on autoencoder (AE) neural networks to compress the datasets and perform the search in an interpretable, compressed latent space. A distance-regularized Siamese twin autoencoder (DIRESA) architecture is designed to preserve distance in latent space while capturing the nonlinearities in the datasets. Using conceptual climate models of different complexities, we show that the latent components thus obtained provide physical insight into the dominant modes of variability in the system. Compressing datasets with DIRESA reduces the online storage and keeps the latent components uncorrelated, while the distance (ordering) preservation and reconstruction fidelity robustly outperform Principal Component Analysis (PCA) and other dimension reduction techniques such as UMAP or variational autoencoders.
Significance Statement
This paper introduces a novel, deep learning based dimension reduction method called DIRESA. We believe that this method has the potential to complement the widely used Principal Component Analysis. It gives the user insight into the dominant modes of variability in the system while offering excellent reconstruction fidelity and distance preservation in latent space.
To increase the impact of this method and facilitate its adoption by the scientific community, we have made it available in a user-friendly Python package. The DIRESA method could thus become a standard tool for many atmospheric and climate scientists.
@@ 1 Introduction
Today, weather and climate models and observation systems generate unprecedented volumes of data. For example, the Coupled Model Intercomparison Project phase 6 (CMIP6) represents more than 20 PB of climate model output data (Li et al., 2023), and the world’s Earth Observation data collections grow by around 100 PB per year (Wilkinson et al., 2024).
@@
These unwieldy archives of Earth system data provide enormous opportunities as well as challenges. They are usually stored in formats such as NetCDF, which has natively supported zlib-based compression since version 4.0. This compression technique does not consider the specific context of climate data, and as a result, many more bits of information are stored than are meaningful or necessary (Klöwer et al., 2021). Classical image or video compression techniques may retain some interpretability but are optimized for specific targets, such as the human visual system. The compressed representation learned by neural image compression methods is more adaptive but generally is not interpretable, nor does it allow for similarity search in the compressed domain (Mishra et al., 2022). Compression can also be done by overfitting a coordinate-based neural network and taking the resulting parameters as a compact representation of the original grid-based data (Huang and Hoefler, 2022), but such an approach still lacks interpretability and distance preservation.
@@
To allow for the full potential of this vast treasure of data to be unlocked, an approach is needed to reduce these datasets in a domain-aware and physically meaningful way. The approach presented in this work uses a Siamese twin AE with regularization terms to anchor the latent space to meaningful properties such as distance (ordering) preservation and statistical independence. The first property is crucial for using algorithms based on distances, such as $k$-means and analogs, in the reduced latent space. The second property reduces redundancies between components and promotes orthogonality in the latent space, ensuring that the latent components represent distinct and independent features with a higher potential for interpretability.
@@
A major application that benefits from long records of reanalysis and climate model data is the search for analogs. Methods based on analogs, which rely on the retrieval of similar weather patterns, are used for forecasting (Van den Dool, 1989), as Lorenz already proposed in 1969 (Lorenz, 1969), or for the calculation of local attractor properties, such as the instantaneous dimension, a proxy for predictability (Faranda et al., 2017). In analog data assimilation, the analogs can be searched in historical data archives (Lguensat et al., 2017) or constructed via generative machine learning methods, such as variational AE (Yang and Grooms, 2021) or generative adversarial networks. Besides statistical and dynamical downscaling, higher resolution weather and climate information can be obtained using analog methods (Zorita and Von Storch, 1999; Ghilain et al., 2022) or neural network-based techniques (Rozoff and Alessandrini, 2022). Finally, analogs are a cornerstone of a novel climate change attribution approach that accounts for changes in atmospheric circulation patterns (Faranda et al., 2022).
@@
As a downside to the growing datasets, however, looking up analogs in these vast records becomes increasingly computationally expensive as the dimensionality of the data and length of the datasets increase. As looking up analogs by minimizing Euclidian distance can be practically infeasible, many approaches have been proposed to quickly retrieve similar patterns of one or more atmospheric variables. These are based on dimensionality reduction (DR) techniques such as the popular classical PCA method. To capture the nonlinearity of the data, other techniques were developed (Van Der Maaten et al., 2009). UMAP (Uniform Manifold Approximation and Projection) is one such technique (McInnes et al., 2018), which can be combined with advanced search algorithms to find the analog in a fast and accurate way (Franch et al., 2019). DR is also combined with other data analysis techniques, such as clustering (Neal et al., 2016; Chattopadhyay et al., 2020), or topological techniques, such as persistent homology (Strommen et al., 2022), for finding weather regimes.
@@
Autoencoders are another technique, based on deep learning, to produce a lower-dimensional latent representation for a dataset or distribution (Manning-Dahan, 2018). However, they lack the statistical independence of the latent space components and the ordering of these components. The latter can be overcome by adding a MaskLayer (Royen et al., 2021). The correlation between the components is minimized by adding a covariance regularization term on the latent space layer (Ladjal et al., 2019).
@@
When using autoencoders, the Euclidian distance between two samples in the original space is not necessarily correlated with the distance in latent space. Our architecture uses a Siamese twin AE (see Figure 1) to preserve distance (ordering) in latent space. The two encoder branches share the weights and produce two different latent space representations. These are used for the distance loss term, similar to the invariance term used by Bardes et al. (2021), which ensures that the distance between samples in the dataset is reflected (preserved or correlated) in the distance between latent representations of those samples. The Siamese twin architecture is also used in AtmoDist for the task of predicting temporal distance between samples to obtain an informative latent representation of the atmospheric state (Hoffmann and Lessig, 2023).
@@ 2.1 Lorenz ’63
The Lorenz ’63 system (Lorenz, 1963) is a 3-dimensional simplified model for atmospheric convection. This system of three ordinary differential equations (ODE) displays chaos and features a strange attractor with the famous butterfly-wings shape for specific ranges of its parameters (Palmer, 1993). The 3-dimensional Lorenz butterfly is compressed into a 2-dimensional latent space by the different DR techniques. The system of ODEs with parameters $\sigma=10$, $r=28$, and $b=8/3$ was integrated using the 4th-order Runge-Kutta numerical scheme, from starting point (1,0,1) with a timestep of 0.0025. The first 1000 points are ignored as they are considered part of the transient. The next 100,000 steps are split into a training dataset (first 80,000), a validation (next 10,000), and a test dataset (last 10,000). Before feeding the data into the DR methods, it is scaled to $[0,1]$. Important to remark is that the Hausdorff fractal dimension of the Lorenz butterfly is about 2.06, so it is theoretically impossible to map the 3-dimensional dataset into two dimensions while retaining all the information.
@@ 2.2 Modular Arbitrary-Order Ocean-Atmosphere Model
MAOOAM (Modular Arbitrary-Order Ocean-Atmosphere Model) is a quasigeostrophic coupled ocean-atmosphere model for midlatitudes (De Cruz et al., 2016; Vannitsem et al., 2015). It features a 2-layer atmosphere that is mechanically and thermodynamically coupled to a shallow-water ocean layer in the $\beta$-plane (linearized Coriolis) approximation. The DDV2016 setup was used, with an atmospheric and oceanic resolution of 2x–2y and 2x–4x, resulting in a 36-dimensional model, with ten wavenumbers for the atmospheric stream function and temperature and eight for the ocean stream function and temperature. The dataset was obtained by integrating the model with the 2nd-order Runge-Kutta numerical scheme (with nondimensional timestep 0.01) for $4\times 10^{8}$ timesteps, with a write-out every 100th timestep (1 timestep is $1/f_{0}$ s, with $f_{0}$ the Coriolis parameter at 45°N, which is about 2.7 hours). The first 1 million points ($10^{8}$ timesteps) are ignored as part of the transient, required due to the slow oceanic dynamics. For the following 3 million points, which represent 92,000 years of data, the wavenumbers are converted to a grid of 8 x 6 and four channels (for the four variables) and then split up into a training (first 2.4 million), a validation (next 0.3 million), and a test dataset (last 0.3 million).
@@ 3.1 PCA and KPCA
PCA linearly transforms the data into new coordinates, the principal components (Pearson, 1901; Hotelling, 1933). These components are the eigenvectors of the covariance matrix of the dataset. The amount of dataset variation in the direction of each eigenvector determines the ordering of the coordinates. As the covariance matrix is symmetric, the eigenvalues are real, and the eigenvectors are orthogonal. For dimensional reduction, the coefficients of the first principal components can be used as latent representation, as they explain most of the variance in the dataset.
In KPCA, the traditional PCA is combined with kernel methods (Schölkopf et al., 1997). The original data points are mapped non-linearly on a high dimensional Hilbert space, where the PCA is applied. The so-called kernel trick is used to avoid working in this higher dimension (and so being highly computationally expensive). The calculations are done in the original space, replacing the standard inner product with a kernel function representing the inner product in the Hilbert space.
@@ 3.2 UMAP
UMAP is based on manifold learning and topological data analysis (McInnes et al., 2018). It creates a fuzzy topological representation of the data, a weighted graph, by connecting the points within a certain radius, which is calculated locally based on the distance to the nearest neighbors. A stochastic gradient descent algorithm with a cross-entropy loss function searches for a low-dimensional representation with the closest possible similar fuzzy topological structure.
@@ 3.3 Autoencoders
An AE is a deep-learning artificial neural network (ANN) that consists of two parts: an encoder that compresses the data into a latent representation and a decoder that decompresses this representation into its original form (Kramer, 1991). The AE is trained by minimizing the reconstruction loss, which ensures that the output of the decoder is as similar as possible to the encoder’s input. We use the MSE between the encoder’s input and the decoder’s output as reconstruction loss. After the training and testing of the neural network are done, the original dataset can be compressed with the encoder.
@@
In a VAE, the latent components are not single numbers but probability distributions, usually Gaussian distributions (Kingma and Welling, 2022). The encoder has three output layers: one for the means, one for the variances (or logarithm of the variances) of the distributions, and one random sampling taken from those distributions that feed into the decoder submodel. A VAE is a generative method, as each sample of the latent distributions produces a different output after decoding. A VAE has two loss functions. The first one is the reconstruction loss, as in a standard AE. The second loss function contains a Kullback–Leibler divergence regularization term (KL loss), which forces the latent distributions to be standard normal distributions. The total loss function is the weighted average of both losses.
Figure 2: (a) Requirement for the AE: Reconstructed data as close as possible to the original. (b) DR requirements: Distance in original space is related to distance in latent space, and latent components are independent and ordered in terms of importance.
@@ 3.4 DIRESA
    Reconstructed dataset as close as possible to the original dataset;
@@
    The distance between points in the original space is reflected in their distance in latent space;
@@
    The latent components are independent of each other;
@@
    The latent components are ordered in terms of importance.
The last three requirements focus on capturing the essential structure and relationships within the data in a lower-dimensional space, while the first one is specific to autoencoders. The DIRESA architecture (Figure 1) fulfils the above requirements. The original dataset $\mathbf{X}=\left(\mathbf{x}_{i}\right)_{i=1\ldots n}$, with $n$ the number of data points, is encoded by the encoder $f_{\theta}$, parametrized by $\theta$, into a latent representation $\mathbf{Z}$
$f_{\theta}(\mathbf{X})=\mathbf{Z}\in\mathbb{R}^{n\times L}$
where $L$ is the number of latent components. The decoder $g_{\phi}$, parametrized by $\phi$, decodes the latent dataset back into the original space
$g_{\phi}(\mathbf{Z})=\mathbf{Y}.$
@@
As in a standard AE, the reconstruction loss enforces the first requirement. We use the MSE between the original data, fed into the encoder, and the decoder output as reconstruction loss
$\mathcal{L}^{MSE}_{recon}=\frac{1}{n}\times\sum_{i=1}^{n}\left\|\mathbf{x}_{i}-\mathbf{y}_{i}\right\|^{2}.$
The reconstruction loss is related to the fraction of variance unexplained $FVU=1-R^{2}$, with $R^{2}$ being the coefficient of determination
$R^{2}=1-\frac{\sum_{i=1}^{n}\left\|\mathbf{x}_{i}-\mathbf{y}_{i}\right\|^{2}}{\sum_{i=1}^{n}\left\|\mathbf{x}_{i}-\mathbf{\bar{x}}\right\|^{2}}.$
In the case of PCA, $R^{2}$ corresponds to the percentage of explained variance and the FVU is equal to the unexplained variance.
@@
The twin encoder shares the weights and biases with the original encoder but receives other samples from the data set $\mathbf{X}^{\prime}=\left(\mathbf{x}^{\prime}_{j}\right)_{j=1,\ldots,n}$ , which are obtained by shuffling the order of the data. The twin encodes a data point $\mathbf{x}^{\prime}$ into $\mathbf{z}^{\prime}$. The random ordering of the shuffled dataset can be done beforehand on the complete dataset (and so does not change during training) or can be executed per batch during training by the data generator. We used the first option here to be independent of the batch size. A distance layer calculates the Euclidian distance $\left\|\mathbf{x}-\mathbf{x^{\prime}}\right\|$ between the original and shuffled input and between the latent components of the two twins $\left\|\mathbf{z}-\mathbf{z^{\prime}}\right\|$. Minimizing the distance loss forces the relationship between those two. For DIRESA, several distance loss functions have been implemented. $\mathcal{L}^{MSE}_{dist}$ uses the MSE
$\mathcal{L}^{MSE}_{dist}=\frac{1}{n}\times\sum_{i=1}^{n}(\left\|\mathbf{x}_{i}-\mathbf{x^{\prime}}_{i}\right\|-\left\|\mathbf{z}_{i}-\mathbf{z^{\prime}}_{i}\right\|)^{2}$
meaning that distance is preserved between the original and latent space. The MSLE, which calculates the MSE between $log(\left\|\mathbf{x}-\mathbf{x^{\prime}}\right\|+1)$ and $log(\left\|\mathbf{z}-\mathbf{z^{\prime}}\right\|+1)$, is used for $\mathcal{L}^{MSLE}_{dist}$, focusing more on short than long distances. $\mathcal{L}^{Corr}_{dist}$ uses a distance Pearson correlation loss
$\mathcal{L}^{Corr}_{dist}=1-Corr(\left\|\mathbf{X}-\mathbf{X^{\prime}}\right\|,\left\|\mathbf{Z}-\mathbf{Z^{\prime}}\right\|)$
forcing the correlation between original and latent space distances. The distance Pearson correlation loss is calculated per batch (one figure per batch and not one figure per sample as in an MSE loss function). Therefore, the batch size must be large enough to provide a reliable estimate of the correlation. A last distance loss function is $\mathcal{L}^{LogCorr}_{dist}$, where the correlation loss is computed on the logarithm of the distances plus one.
The independence of the latent components is forced by the covariance loss
$\mathcal{L}_{cov}=\frac{1}{L\times(L-1)}\times\sum_{i\neq j}cov^{2}_{i,j}(\mathbf{Z})$
@@
which is the normalized squared sum of the non-diagonal terms of the covariance matrix. The denominator keeps the loss function independent of the latent space dimension. The covariance matrix is calculated by batch, so again, it is essential to have the batch size big enough so that the batch covariance is a good approximation of the dataset covariance.
@@
The total loss is a weighted average of the three different loss components. For lowering the weight factor tuning effort, annealing (Behrens et al., 2022) is used for the covariance loss, meaning that the weight factor is 0 at the start and gradually increases. Annealing stops when the covariance loss reaches a point, where we consider the latent components independent. The covariance regularized AE (CRAE) (Ladjal et al., 2019) is similar to DIRESA but with only the covariance loss and no distance loss (and so no twin encoder).
@@
The last requirement, the ordering of the latent components, could be imposed by a Masklayer, but in practice, it was found to hamper the training of the neural network. Therefore, the ordering is not imposed during training but is calculated afterwards, based on the $R^{2}$ score. The procedure for this is described in Appendix E.
In Appendix E, an ablation study demonstrates the usefulness of the different loss functions. The ablation study shows that the reconstruction loss can be left out, which yields a simplified DIRESA architecture. However, without the reconstruction loss, there is no decoder for the latent space dataset, and without a decoder, the ordering of the components, according to the procedure in Appendix D, cannot be calculated.
@@
Figure 3: Lorenz test dataset (blue) next to the decoded random uniform latent components (orange) for the different DR techniques
Figure 4: Encoded Lorenz test dataset for the different DR techniques